### v1

# Protecting Deep Learning Model Copyrights with Adversarial Example-Free Reuse Detection

Abstract
Model reuse techniques can reduce the resource requirements for training high-performance deep neural networks (DNNs) by leveraging existing models. However, unauthorized reuse and replication of DNNs can lead to copyright infringement and economic loss to the model owner. This underscores the need to analyze the reuse relation between DNNs and develop copyright protection techniques to safeguard intellectual property rights. Existing DNN copyright protection approaches suffer from several inherent limitations hindering their effectiveness in practical scenarios. For instance, existing white-box testing-based approaches cannot address the common heterogeneous reuse case where the model architecture is changed, and DNN fingerprinting approaches heavily rely on generating adversarial examples with good transferability, which is known to be challenging in the black-box setting. To bridge the gap, we propose NFARD, a Neuron Functionality Analysis-based Reuse Detector, which only requires normal test samples to detect reuse relations by measuring the models’ differences on a newly proposed model characterization, i.e., neuron functionality (NF). A set of NF-based distance metrics is designed to make NFARD applicable to both white-box and black-box settings. Moreover, we devise a linear transformation method to handle heterogeneous reuse cases by constructing the optimal projection matrix for dimension consistency, significantly extending the application scope of NFARD. To the best of our knowledge, this is the first adversarial example-free method that exploits neuron functionality for DNN copyright protection. As a side contribution, we constructed a reuse detection benchmark named Reuse Zoo that covers various practical reuse techniques and popular datasets. Extensive evaluations on this comprehensive benchmark show that NFARD achieves F1 scores of 0.984 and 1.0 for detecting reuse relationships in black-box and white-box settings, respectively, while generating test suites $2\sim 99$ times faster than previous methods.
Index Terms:
copyright protection, reuse detection, neuron functionality, deep neural networks
I Introduction
Deep learning has been extremely successful in a wide range of tasks [1, 2, 3]. However, training a high-performance deep neural network (DNN) requires not only a large amount of labeled data but also significant computational resources that escalate with the complexity of the model. For example, it cost more than \$100 million to train GPT-4 [4], one of the most advanced DNNs in existance. Meanwhile, just like traditional software, deep learning models are often reused and modified for another learning task to save resources. To this end, many model reuse techniques have been proposed to avoid training models from scratch. For example, transfer learning [5] reduces the required data by applying knowledge gained on one problem to another different but similar task. Model compression [6] reduces the model size while preserving reasonable performance so that DNNs can be deployed in resource-constrained scenarios.
Despite the benefits, model reuse techniques also pose concerns about copyright infringement. DNN models can be exposed in multiple ways without adequate protection, including distributed copies and Machine Learning as a Service (MLaaS) [7]. Unauthorized reuse, such as commercialization in breach of license or theft via model extraction techniques [8, 9, 10] (a form of model reuse), can lead to copyright infringement, posing threats to the competitive edge of the DL model owner and causing significant economic loss. With the push for AI governance, such as the OECD recommendation on AI [11] and the European Union’s Artificial Intelligence Act [12], DNN copyright protection is becoming increasingly important to the responsible development of deep learning models in both academic and industrial practices. There is a growing need to develop DNN copyright protection methods to prevent the valuable intellectual properties from being illegally replicated and to provide the necessary evidence to resolve copyright infringement disputes.
Several types of DNN copyright protection methods have been proposed, including watermarking, fingerprinting, and testing-based similarity comparison. Watermarking methods [13, 14, 15, 16, 17] embed a secret watermark into the model during training by utilizing over-parameterization. Then the owner can identify whether a suspect model is a copy by comparing the extracted watermarks. However, the watermark needs to be embedded during the training phase, which may compromise the performance and security of the model. Moreover, recent works have shown that adaptive attacks [18, 19] can remove watermarks. Fingerprinting methods [20, 21, 22], on the other hand, do not tamper with the training procedure. Instead, it extracts a fingerprint from the model using a set of test samples, which contains characterizing features of the model that remain unchanged after applying DNN reuse techniques. Any model with a similar fingerprint will be judged as a copy. Most existing fingerprinting methods heavily rely on generating adversarial examples with good transferability, which is known to be challenging, especially in the black-box setting [23]. Another line of testing-based similarity comparison approaches [24, 25] employs a set of testing metrics and specially crafted test samples to measure the models’ similarity to detect reuse relations. However, existing white-box testing methods cannot address the common heterogeneous reuse case where the model architecture is changed. In summary, existing DNN copyright protection approaches suffer from inherent limitations that hinder their practical use in diverse scenarios.
To address the gap, we propose NFARD, a Neuron Functionality Analysis-based Reuse Detector, which only requires normal test samples to detect reuse relations by measuring the models’ differences on a newly proposed model characterization, i.e., neuron functionality (NF). Neuron functionality has been well practiced to characterize a DNN in neural network representation studies [26, 27]. In this work, we leverage neuron functionality to protect DNN copyright. Specifically, a neuron’s functionality is the mapping from the input space to the one-dimensional output space of the neuron, i.e., $\mathcal{X}\rightarrow\mathbb{R}$ with $\mathcal{X}$ representing the input space. [^1] Neuron functionality is the direct mapping function which is fundamentally different from neuron output statistics used in previous approaches. Recent studies [26, 27, 28] have shown that independently trained neural networks have divergent neuron functionalities. On the other hand, neural networks obtained by reuse techniques have similar neuron functionalities to those of the original model. Therefore, we can detect model reuse relations by measuring the similarity of neuron functionalities, which can be approximated only using normal samples, avoiding the trouble of carefully generating fingerprints or test samples. We devise a set of distance metrics to measure the similarity of neuron functionality under both white-box and black-box settings. Besides, in order to extend the application scope of NFARD, we further propose a linear transformation method to map the neuron functionality of neural networks with different architectures or classification tasks to a common feature dimension. The key is to identify the optimal projection matrix from high-dimensional features to a low-dimensional model representation. We compute the optimal projection by solving a set of least squares problems. This enables NFARD to effectively address heterogeneous cases, facilitating its utility in diverse practical scenarios.
To enable a comprehensive evaluation of NFARD, we construct a model reuse detection benchmark named Reuse Zoo, which consists of 250 neural network models, covering seven commonly used model reuse techniques, three datasets, and four popular model architectures. We use this benchmark to evaluate NFARD and compare it with three baseline approaches: a representative fingerprinting method, IPGuard [20], and two state-of-the-art testing-based similarity comparison approaches, i.e., ModelDiff [25] and DeepJudge [24]. Extensive experimental results show that NFARD can effectively detect reuse relations in both black-box and white-box settings, achieving an F1-score of 0.984 and 1.0, respectively, and is $2\sim 99$ times faster than previous methods in terms of test suite generation.
In summary, our main contributions are as follows:
  1. 1.
    We propose NFARD, a neuron functionality analysis-based reuse detector for DNN copyright protection, which incorporates a new model characterization and a set of distance metrics in both white-box and black-box settings. Compared to existing approaches, NFARD does not rely on generating fingerprints or test samples.
  2. 2.
    We propose a linear transformation approach via least squares solving to address the heterogeneous reuse case, significantly extending the application scope of the proposed reuse detector.
  3. 3.
    We construct and release a model reuse benchmark with 250 models for evaluating DNN copyright protection methods, covering a variety of popular model reuse types.
  4. 4.
    We extensively evaluate NFARD to show that it has a broader application scope and can achieve better detection results than SOTA methods, especially on heterogeneous cases.
II Background
II-A Deep Neural Network and Neuron Functionality
We focus on classification tasks in this work. A deep neural network classifier is a function mapping from the input space to a probability distribution, i.e., $f:\mathcal{X}\to[0,1]^{m}$ where $\mathcal{X}$ is the input space and $m$ is the number of classes. The neural network includes $L$ layers $f^{(1)},f^{(2)},\cdots,f^{(L)}$ whose output dimensions are $m_{1},m_{2},\cdots,m_{L}$, respectively. We assume $f^{(L)}$ is a linear classifier and call it the last layer of the model. Given an input $\bm{x}\in\mathcal{X}$, the output of $f$ on $\bm{x}$ is defined by $f=\sigma\circ f^{(L)}\circ\cdots\circ f^{(1)}$, where $\sigma$ is the softmax function that converts any vector to a probability vector.
In a neural network $f$, the functionality of a neuron is the mapping from the input space to the one-dimensional output space of the neuron, i.e., $f_{i}^{(k)}:\mathcal{X}\to\mathbb{R}$, where $f_{i}^{(k)}$ is the $i$-th neuron of the $k$-th layer. Mathematically, the functionality of a neuron $f_{i}^{(k)}$ is a highly non-convex function. In practice, it is approximated by a set of normal samples $\{\bm{x}_{j}\}_{j=1}^{n}$ and characterized as a vector $\bm{\mathrm{nv}}(k,i)$:
$\bm{\mathrm{nv}}(k,i)=(f_{i}^{(k)}(\bm{x}_{1}),f_{i}^{(k)}(\bm{x}_{2}),\cdots,f_{i}^{(k)}(\bm{x}_{n}))^{\intercal},$ (1)
We refer to the vector form neuron functionality $\bm{\mathrm{nv}}(k,i)$ as the neuron vector [^2] This is also called activation vector in some other work [27], but it is different from layer output on a single sample. in this work. To the extent of our knowledge, the concept of neuron functionality was first formally introduced in [29] to prune neural networks by identifying neurons with similar functionalities.
Recent work [26, 27] on neural network representation proposed a neuron functionality analysis approach to study DNN learning dynamics and interpretability. They have shown that independently trained DNN models with different weight initialization have different corresponding neuron functionalities, which is in accordance with [28]’s conclusions. That is to say, even if two DNN models $f$ and $g$ have the same structure, training dataset, and training setting, their corresponding neuron functionalities $\bm{\mathrm{nv}}_{f}(k,i)$ and $\bm{\mathrm{nv}}_{g}(k,i)$ can vary greatly as long as they have different weight initialization.
II-B Model Reuse Detection for Copyright Protection
We aim to safeguard the copyright of DNNs by detecting model reuse relations. Initially, the model owner trains a victim model using his/her own resources, including the labeled data and the computational resource. Then model reuse methods are applied to the victim model to obtain a surrogate model, or a reused copy, which demonstrates comparable performance at a lower cost on similar or new tasks. We consider the assessment of whether a reuse relationship constitutes an unauthorized exploit as a non-technical issue and focus on the detection of the reuse relation.
Given a victim model $f_{v}$ and a suspect model $f_{s}$, the task of the reuse detector is to determine whether $f_{s}$ is a positive suspect model, i.e., a surrogate model obtained by applying reuse methods to $f_{v}$, or a negative suspect model trained from scratch or irrelevant to $f_{v}$. The reuse detector may require a set of negative suspect models $\{f_{r,i}\}_{i=1}^{R}$ trained independently of the victim model serving as references to make the decision. Such models are called reference models [^3] In practice, hyperparameter tuning process (e.g., grid search) produces multiple independently trained models that can be used as reference models.. Depending on the access of the reuse detector to the models, the detection methods can be divided into two categories: white-box and black-box methods. In the white-box setting, the detector has access to the internals of the neural networks and the final probability output. While in the black-box setting, the detector only has access to the final probability output. In either case, we assume that a subset of the training data of the victim model is available for the reuse detector (e.g., the model owners themselves are searching for surrogate models).
II-C Model Reuse Methods
We consider seven commonly used DNN reuse methods, including fine-tuning, retraining, pruning, quantization, knowledge distillation, transfer learning, and model extraction. These reuse methods are classified into homogeneous and heterogeneous methods depending on whether the obtained surrogate model has the same structure as the victim model.
Fine-tuning and retraining techniques train the victim model for a few epochs with a relatively small learning rate, and some weight parameters may be frozen [30] to enable a finer control. The difference between them is that retraining also re-initializes the weights of the last layer. Pruning and quantization methods are compression techniques that aim to reduce DNN size by removing trivial weights or quantizing parameters to low-bit values while minimizing the drop in model performance [31, 32]. Surrogate models obtained by these four methods have the same model structure as the victim model, so they belong to homogeneous methods. On the other hand, knowledge distillation and model extraction train surrogate models by exploiting the victim model’s outputs [33, 9, 10], and they usually result in a different architecture; transfer learning replaces the last layer with a new one [5], and the classification task becomes different. Therefore, these three reuse techniques change the model structure and fall into the category of heterogeneous methods.
Detecting heterogeneous cases is usually more challenging. To the best of our knowledge, in previous work, only ModelDiff [25] could handle transfer learning reuse cases, but it is not good at dealing with the other two heterogeneous reuse types. We propose a benchmark Reuse Zoo covering different types of reuse methods for DNN reuse detector evaluation, and we elaborate on the construction details in Section IV.
III Methodology
This section presents NFARD, the proposed neuron functionality analysis-based reuse detection method. We first conduct a preliminary study to demonstrate the principle of our approach. Then we present the details of NFARD. After that, we describe how to handle heterogeneous reuse cases by the linear transformation method.
III-A Insight into Reused DNN Models
Training a DNN model involves optimizing the weights to reach a local minimum of the loss function. Due to the non-convexity of neural networks, independently trained models may end up at different local minima with distinct weights. In contrast, reusing a trained model to obtain a surrogate model is essentially finding another set of weights close to the original ones. The more the reuse method leverages the victim model’s knowledge, the more similar the obtained surrogate model to the victim model. Inspired by this observation, we conduct a preliminary experiment to see if the neuron functionality of a reused model is also more similar to that of the victim model compared to independently trained reference models.
To begin with, we introduce two metrics to measure the similarity of neuron functionality. The first metric is $dist_{eu}^{(k)}$, which is the average Euclidean distance between the $k$-th layer’s neuron vectors of two models, defined as:
$dist_{eu}^{(k)}(f_{v},f_{s},X)=\dfrac{1}{m_{k}}\sum\limits_{i=1}^{m_{k}}\lVert\bm{\mathrm{nv}}_{v}(k,i)-\bm{\mathrm{nv}}_{s}(k,i)\rVert_{2},$
where $X=\{\bm{x}_{1},\cdots,\bm{x}_{n}\}$ is a set of test inputs, $\bm{\mathrm{nv}}_{v}(k,i)$ is the $i$-th neuron vector of the $k$-th layer of $f_{v}$, and $\bm{\mathrm{nv}}_{s}(k,i)$ is that of $f_{s}$. The second metric is the average cosine distance of the neuron vectors, denoted as $dist_{ac}^{(k)}$:
$dist_{ac}^{(k)}(f_{v},f_{s},X)$ $=\dfrac{1}{m}\sum\limits_{i=1}^{m}dist_{cos}(\bm{\mathrm{nv}}_{v}(k,i),\bm{\mathrm{nv}}_{s}(k,i)),$
where $dist_{cos}$ represents cosine distance. Compared with $dist_{eu}^{(k)}$, $dist_{ac}^{(k)}$ focuses more on the difference in direction rather than the scale. The smaller the metrics, the more similar the two models are in terms of neuron functionality. We do not include metrics used in previous DNN representation studies (such as SVCCA similarity [26]) because they target analyzing training dynamics on hidden layers where correlation analysis is essential, and our goal is to detect reuse relations using the output layer neuron vectors.
(a) GoogLeNet-CIFAR10
(b) GoogLeNet-CIFAR100
(c) MobileNet V2-CIFAR10
(d) MobileNet V2-CIFAR100
Figure 1: Distance metrics between victim and suspect models.
Next, we examine four victim models with different structures and learning tasks. For each victim model, we prepare 8 surrogate models by applying homogeneous reuse methods and 40 reference models that are independently trained on the same task (details of these models are available in Section IV). We compute the distance metrics $dist_{eu}^{(L)}$ and $dist_{ac}^{(L)}$ on the last layer (before the softmax function) using 1000 normal samples, and the results are presented in Figure 1, where surrogate models and reference models are plotted in orange and blue, respectively. We can see that reference models have larger distance values and are concentrated in the upper right corner of the plane, while surrogate models have smaller distance values and are scattered on the lower left part. Such differences indicate that homogeneous surrogate models are more similar to victim models than independently trained models regarding neuron functionality, which is exactly the principle of our proposed method.
We use the output logits to compute the metrics in this preliminary experiment. But in practice, we can only get probability prediction obtained by applying softmax to logits. Measuring the similarity of neuron functionality in such a black-box setting is one challenge that we need to address, and another challenge is dealing with heterogeneous reuse types. Nevertheless, the preliminary results give us insight into the reused DNN models.
III-B Neuron Functionality Analysis-based Reuse Detector
Based on the finding in our preliminary study, we propose NFARD, a neuron functionality analysis-based reuse detector. We depict the detection procedure in Algorithm 1, which consists of three core components: test suite selection, distance metrics, and decision criteria. We first concentrate on dealing with homogeneous cases in white-box and black-box settings and present the linear transformation method for heterogeneous cases later.
III-B1 Test suite selection
NFARD uses a subset of the victim model’s training data to extract neuron vectors to compute the neuron functionality similarity between different neural networks. Note that NFARD does not require adding any perturbations to the test samples. To induce aligning behavior between the victim and surrogate models, we construct the test suite using samples that elicit the least confident predictions from the victim model, where the degree of confidence is measured by the entropy of the predicted probabilities.
Input: victim model $f_{v}$, suspect model $f_{s}$, test suite size $n$, dataset $D$, reference models $\{f_{r,i}\}_{i=1}^{R}$
Output: decision on if $f_{s}$ is a surrogate of $f_{v}$
$X\leftarrow TestSuiteSelection(D,n)$;
$distances\leftarrow DistanceMetrics(f_{v},f_{s},X)$;
$\{ref_{i}\}_{i=1}^{R}\leftarrow DistanceMetrics(f_{v},\{f_{r,i}\}_{i=1}^{R},X)$;
$decision\leftarrow DecisionCriteria(distances,\{ref_{i}\}_{i=1}^{R})$;
return $decision$
Algorithm 1 Reuse detection procedure of NFARD
III-B2 Distance metrics
NFARD employs different distance metrics for the white-box and black-box cases and allows extending our method by incorporating more metrics.
In the white-box setting, the detector has access to the neuron vectors of all layers, so we can directly adopt $dist_{eu}^{(k)}$ and $dist_{ac}^{(k)}$ as the distance metrics. The parameter $k$ specifies the target layer from which we extract the neuron vectors, and the choice of target layer depends on neural network architecture. In practice, we usually choose a relatively shallow layer as the target layer.
In the black-box setting, neither $dist_{eu}^{(k)}$ nor $dist_{ac}^{(k)}$ is directly applicable as we have the probability predictions (by applying softmax function) rather than the neuron vectors of the last layer. It is impossible to recover the output (neuron vector) of the last layer from the probability predictions since the softmax function is not bijective. We propose to take the logarithm of the probability to compute a logit approximation. The differences in the computation of distance metrics in the black-box setting are shown as follows:
$dist_{eu}^{\prime}(f_{v},f_{s},X)$ $=\dfrac{1}{m}\sum\limits_{i=1}^{m}\lVert\tilde{\bm{\mathrm{nv}}_{v}}(L,i)-\tilde{\bm{\mathrm{nv}}_{s}}(L,i)\rVert_{2},$
$dist_{ac}^{\prime}(f_{v},f_{s},X)$ $=\dfrac{1}{m}\sum\limits_{i=1}^{m}dist_{cos}(\tilde{\bm{\mathrm{nv}}_{v}}(L,i),\tilde{\bm{\mathrm{nv}}_{s}}(L,i)),$
where $\tilde{\bm{\mathrm{nv}}}(L,i)$ is the last layer’s approximated neuron vector ($f(\bm{x})_{i}$ is the $i$-th element of the predicted probability on $\bm{x}$):
$\tilde{\bm{\mathrm{nv}}}(L,i)=(\log f(\bm{x}_{1})_{i},\log f(\bm{x}_{2})_{i},\cdots,\log f(\bm{x}_{n})_{i})^{\intercal}.$
Taking the logarithm of the probability amplifies the differences between models because the probability takes a value from 0 to 1, while its logarithm takes a value from $\infty$- ∞ to 0. We will demonstrate the advantage of the logarithm approximation in Section V.
Note that our metrics are defined on the column vectors of the neuron matrices, i.e., neuron vectors, while the metrics of existing methods are mostly defined on the layer vectors, usually calculated by running a single sample and collecting layer outputs.
III-B3 Decision criteria
NFARD requires a set of reference models to make the final decision. For each distance metric $dist_{i}$, we define its decision function as:
$d_{i}(x,y_{1},\cdots,y_{s})=\tilde{z}-\alpha\cdot\textrm{IQR}-x,$
where $x$ is the distance between $f_{v}$ and $f_{s}$, $y_{i}$ is the distance between $f_{v}$ and $f_{r,i}$, $\tilde{z}$ and IQR are the median and the interquartile range of $\{x,y_{1},\cdots,y_{s}\}$, and $\alpha$ is a pre-fixed parameter. Conventionally, data points below $\tilde{z}-1.5\ \textrm{IQR}$ are regarded as outliers (known as the IQR rule). In our method, we replace the coefficient 1.5 with a user-specified parameter to control the trade-off between sensitivity and specificity. The higher the decision function $d_{i}$ value, the greater the deviation of the suspect model from the reference models, indicating a higher likelihood of it being a surrogate model.
Considering the varying significance and scales of different distance metrics (e.g., cosine distance is bounded, but Euclidean distance is not), the final decision is determined based on the value of the weighted sum
$decision(x,y_{1},\cdots,y_{s})=\sum_{i=1}^{p}w_{i}\cdot d_{i}(x,y_{1},\cdots,y_{s}),$
where $w_{i}$ is the weight of the distance metric $dist_{i}$. If the weighted sum is greater than 0, NFARD classifies the suspect model as a true surrogate model. Otherwise, the suspect model is considered an independently trained model.
III-C Handling Heterogeneous Reuse Cases
Figure 2: Illustrative example for linear transformation method.
The detection workflow described above only applies to homogeneous cases where the victim and suspect models have the same classification task (in the black-box setting) or the same architecture (in the white-box setting). However, there are also heterogeneous cases that we need to address.
Without loss of generality, suppose we have two (approximated) neuron matrices $\bm{H}_{1}\in\mathbb{R}^{n\times a}$, $\bm{H}_{2}\in\mathbb{R}^{n\times b}$ and $a\geq b$, where the (approximated) neuron matrix $\bm{H}$ is defined as:
$\bm{H}=\begin{pmatrix}\bm{\mathrm{nv}}(k,1)&\bm{\mathrm{nv}}(k,2)&\cdots&\bm{\mathrm{nv}}(k,m_{k})\end{pmatrix}.$
The main obstacle here is that we cannot directly compare the neuron vectors of $\bm{H}_{1}$ and $\bm{H}_{2}$ because there is no apparent functionality correspondence between the neurons of the two models with different structures. To tackle this problem, we construct a linear transformation $\bm{P}\in\mathbb{R}^{b\times a}$ to project neuron vectors in $\bm{H}_{1}$ into a space of the same dimension as $\bm{H}_{2}$ so that the metrics we designed above become applicable to $\bm{H}_{1}\bm{P}^{\intercal}$ and $\bm{H}_{2}$. The linear transformation $\bm{P}$ should preserve the difference among independently trained models in terms of neuron functionality, as well as the similarity between victim models and surrogate models, to enable reuse relation detection between heterogeneous models.
We construct the desired linear transformation $\bm{P}$ by minimizing the distance between neuron vectors in $\bm{H}_{1}$ and $\bm{H}_{2}$ under a linear relation, which is formulated as the following optimization problem:
$\bm{P}=\operatorname*{argmin}_{\bm{W}}\;\lVert\bm{H}_{1}\bm{W}^{\intercal}-\bm{H}_{2}\rVert_{F},$ (2)
where $\lVert\cdot\rVert_{F}$ is the Frobenius norm. Intuitively, neuron vectors of reference models are largely linearly independent of victim model’s neuron vectors as they are independently trained, so the distances between victim models and reference models will still remain relatively large after the linear transformation.
An illustrative example is provided in Figure 2. We use colored circles to represent 3-dimensional neuron vectors. The victim model is blue, and it has two neuron vectors. In the heterogeneous case, three surrogate models are depicted in red, each with a single neuron vector. Additionally, three independently trained models are colored in green, each also having one neuron vector. The light blue plane represents the subspace spanned by the victim model’s neuron vectors, and our goal is to find a vector in this plane (i.e., a linear transformation from a 2-dimensional subspace to a 1-dimensional subspace) that yields a minimum distance to a given suspect model. This example presents the simplest case where the suspect model has only one neuron vector, and the solution to problem (2) can be obtained by an orthographic projection.
As shown in this example, reference models’ neuron vectors are spread further from the subspace than surrogate models’ due to their lower linear correlation. Therefore, after applying the linear transformation to the larger neuron matrix $\bm{H}_{1}$, we can still distinguish between independently trained models and surrogate models. In fact, as we will show in Section V, distance metrics between surrogate models and victim models after applying the linear transformation method are usually about $10^{-6}\sim 10^{-5}$, while those between reference models and victim models are $1\sim 20$.
With a closer look at problem (2), it solves $b$ sub-problems simultaneously, and each sub-problem can be formulated as
$\bm{p}_{i}=\operatorname*{argmin}_{\bm{w}_{i}}\lVert\bm{H}_{1}\bm{w}_{i}-\bm{\mathrm{nv}}_{2}(k,i)\rVert_{2},$
where $\bm{w}_{i}$ is an $a$-dimensional column vector and $\bm{p}_{i}$ is the $i$-th row vector of $\bm{P}$. Each sub-problem is a standard least square problem, and the closed-form solution of problem (2) is
$\bm{P}=(\bm{H}_{1}^{\dagger}\bm{H}_{2})^{\intercal},$ (3)
where $(\cdot)^{\dagger}$ is the Moore-Penrose inverse, and the proof can be found in many convex optimization textbooks such as [34].
The linear transformation method allows us to apply all metrics above to heterogeneous cases. Specifically, for heterogeneous cases in the black-box case, the transformation is applied after taking the logarithm of the probability. As for heterogeneous cases in the white-box case, we use the second last layer as the target layer. This is because a DNN can be seen as a feature extractor and a linear classifier, and the output of the second last layer is generally the extracted feature vector with low dimension. To make the comparison fair, we apply the linear transformation method to both the suspect and the reference models when the suspect is a heterogeneous model.
IV Benchmark
TABLE I: Complete list of models in the Reuse Zoo benchmark.
Method Configuration Number Examples
Pre-trained 4 pretrain(resnet18,imagenet)
Transfer learning Dataset: CIFAR10, CIFAR100, SVHN 12 transfer(cifar10,resnet50)
Pruning Pruning ratio: 0.3, 0.6 24 transfer(cifar100,mobilenet_v2)-pruning(0.6)
Quantization Data type: 8-bit integers, 16-bit floating points 24 transfer(svhn,googlenet)-quantization(qint8)
Fine-tuning Fine-tune last layer, fine-tune all layers 24 transfer(cifar10,resnet18)-finetune(last)
Retraining Retrain last layer, retrain all layers 24 transfer(cifar10,googlenet)-retrain(last)
Model extraction Architecture: ResNet-18, MobileNet V2, GoogLeNet 9 transfer(cifar100,resnet50)-stealing(resnet18)
Knowledge distillation Architecture: ResNet-18, MobileNet V2, GoogLeNet 9 transfer(svhn,resnet50)-kd(googlenet)
Reference Weight initialization 120 train(resnet18,cifar10,1)
In order to evaluate our approach comprehensively, we create a benchmark named Reuse Zoo consisting of 250 models for model reuse detection. This benchmark contains 4 pre-trained models, 126 surrogate models, and 120 reference models with four popular classification model architectures, including ResNet-18 [35], ResNet-50 [35], MobileNet V2 [36], and GoogLeNet [37].
We collect four models from TorchVision [38] that are pre-trained on ImageNet. For each pre-trained model, we use the transfer learning method to obtain three surrogate models trained on CIFAR10 [39], CIFAR100 [39], and SVHN [40], respectively. Each model is created by replacing the last layer with a new one and then training on the new dataset, which is the most common practice. These 12 models constitute the transfer learning surrogate set whose source victim models are the pre-trained models. We construct eight surrogate models for each transfer learning reuse model by applying pruning, quantization, fine-tuning, and retraining reuse methods. Details of the construction are as follows.
  1. 1.
    Fine-tuning. We adopt two popular fine-tuning strategies: fine-tuning only the last layer and fine-tuning all layers. The fine-tuning surrogate set includes 24 models.
  2. 2.
    Retraining. We first re-initialize the weights of the last layer and then fine-tune it. Similar to the fine-tuning method, we retrain only the last layer and all layers to construct the retraining surrogate set with 24 models.
  3. 3.
    Pruning. We first apply the global pruning method [41] to the victim model with the pruning ratio set as 0.3 and 0.6, respectively, and then fine-tune the pruned models. The pruning surrogate set includes 24 models.
  4. 4.
    Quantization. We use 8-bit integers (qint8) and 16-bit floating points (float16) for quantization. The quantization surrogate set consists of 24 models.
Knowledge distillation and model extraction are typically used to train smaller surrogates with different architectures from the source model. Since the capacity of ResNet-50 is significantly larger than that of the other three architectures, we train three knowledge distillation surrogates and three model extraction surrogates (with ResNet-18, MobileNet V2, and GoogLeNet architectures) for each ResNet-50 transfer learning model. We use the vanilla method [33] for knowledge distillation and the Knockoff method [9] for model extraction. These models constitute the knowledge distillation surrogate set with 9 models and the model extraction surrogate set with 9 models. In total, we obtain a total of 126 surrogate models using these seven model reuse techniques.
We construct 120 reference models in the following way. We independently train ten models from scratch using different weight initialization for each model architecture (four architectures) and each dataset (three datasets). All surrogate and reference models are well-trained and have qualified performance. The complete list of models in the Reuse Zoo benchmark is shown in Table I, with several examples.
This benchmark covers four popular model architectures, ranging from small networks like MobileNet V2 to large networks like ResNet-50, and various reuse methods, notably including heterogeneous reuse methods that can change the model structure. We note that the authors of ModelDiff [25] have proposed a similar DNN reuse detection benchmark referred to as ModelReuse. Compared with ModelReuse, Reuse Zoo (1) is constructed with three commonly used classification datasets while ModelReuse uses two fine-grained image classification (FGIC) datasets that are less representative; (2) includes two more model architectures than ModelReuse, namely ResNet-50 and GoogLeNet, where ResNet-50 is a significantly larger model; (3) covers more reuse techniques, such as fine-tuning and retraining on different layers, which are absent from ModelReuse; (4) is larger, more balanced (126 surrogate models and 120 reference models vs. 84 and 28), and of higher quality (5% accuracy gap between homogeneous models vs. 25% in ModelReuse).
TABLE II: Evaluation results of NFARD under black-box and white-box settings.
Black-box setting White-box setting
Reuse type Detected Decision $dist_{eu}^{\prime}$ $dist_{ac}^{\prime}$ Detected Decision $dist_{eu}^{(k)}$ $dist_{ac}^{(k)}$
Fine-tuning - last 12/12 49.59$\pm$7.95 3.50$\pm$0.79 0.000$\pm$0.000 12/12 154.68$\pm$53.3 0.00$\pm$0.00 0.000$\pm$0.000
Fine-tuning - all 11/12 20.98$\pm$13.3 29.6$\pm$9.20 0.022$\pm$0.014 12/12 150.40$\pm$54.4 3.46$\pm$1.33 0.007$\pm$0.005
Retraining - last 12/12 29.76$\pm$5.70 21.9$\pm$8.49 0.013$\pm$0.003 12/12 154.68$\pm$53.3 0.00$\pm$0.00 0.000$\pm$0.000
Retraining - all 12/12 10.58$\pm$4.93 38.3$\pm$6.48 0.035$\pm$0.011 12/12 146.25$\pm$56.2 6.15$\pm$3.00 0.019$\pm$0.013
Pruning - 0.3 12/12 29.62$\pm$10.9 21.6$\pm$4.30 0.015$\pm$0.010 12/12 149.51$\pm$54.4 4.12$\pm$1.49 0.009$\pm$0.003
Pruning - 0.6 12/12 24.63$\pm$9.95 25.9$\pm$4.31 0.021$\pm$0.010 12/12 142.85$\pm$55.5 8.13$\pm$3.39 0.031$\pm$0.014
Quantization - float16 12/12 52.99$\pm$8.24 0.11$\pm$0.06 0.000$\pm$0.000 12/12 141.86$\pm$62.8 4.98$\pm$8.67 0.065$\pm$0.113
Quantization - qint8 12/12 52.01$\pm$8.16 1.09$\pm$0.32 0.000$\pm$0.000 12/12 141.86$\pm$62.8 4.98$\pm$8.67 0.065$\pm$0.113
Transfer learning 10/12 3.618$\pm$4.06 12.2$\pm$4.44 0.006$\pm$0.005 12/12 142.96$\pm$83.2 16.9$\pm$12.2 0.041$\pm$0.037
Knowledge distillation 9/9 22.92$\pm$8.49 29.0$\pm$6.87 0.025$\pm$0.006 9/9 84.001$\pm$10.6 0.00$\pm$0.00 0.002$\pm$0.004
Model extraction 8/9 12.49$\pm$10.3 37.9$\pm$13.4 0.038$\pm$0.009 9/9 69.243$\pm$20.9 0.00$\pm$0.00 0.125$\pm$0.151
Positive 122/126 126/126
Negative 0/60 -21.74$\pm$9.20 69.0$\pm$12.8 0.056$\pm$0.017 0/60 -68.621$\pm$35.1 129.29$\pm$66.76 0.956$\pm$0.124
V Empirical Evaluation
In this section, we evaluate NFARD in order to answer the following research questions:
  • RQ1: How effective is NFARD at detecting model reuse relation, especially heterogeneous reuse cases?
  • RQ2: Does NFARD have a broader application scope than existing DNN reuse detection methods?
  • RQ3: How does each component affect and contribute to the performance of NFARD?
Parameter setting. We implement NFARD using PyTorch [42] and release its source code along with the benchmark. [^4] https://figshare.com/s/691d5af98dc715cf7009 for anonymous review NFARD has four parameters: the test suite size $n$, the user-specified parameter $\alpha$ for decision criteria, the distance metrics weights $\{w_{i}\}$, and the choice of target layer under the white-box setting. If not specified below, we use a test suite size of $n=1000$, and we empirically set the weights of Euclidean and cosine distances as 1 and 120, respectively. $\alpha$ is set to 0.85 for the black-box case and 3.5 for the white-box case. The target layer is set to the layer at about 25% depth. We will discuss the choice of $\alpha$ and target layer later.
Evaluation method. We use the Reuse Zoo benchmark to evaluate the performance of reuse detectors. The independently trained models in the benchmark are evenly divided into two folds: the first fold serves as reference models for decision-making, and the second fold is used as a negative suspect group to evaluate the false positive rate of the reuse detectors. We employ five reference models with the same architecture and dataset as the victim model to detect whether a suspect model is a true surrogate. One exception is that when the victim model is pre-trained, we do not have enough models trained on ImageNet to serve as reference models (due to limited resources). In this case, we use five reference models with the same architecture and dataset as the surrogate model.
V-A RQ1: Performance of NFARD
V-A1 Benchmark evaluation result
We evaluate the effectiveness of NFARD in reuse relation detection against various reuse methods under black-box and white-box settings. The results are shown in Table II, which includes the number of suspects classified as surrogate models, the average decision values, and the mean and standard deviation of distance metrics. The larger the decision value, the more likely the suspect is a surrogate model. In contrast, the larger the distance, the less likely the suspect is a surrogate model.
According to the results, NFARD achieves 100% precision and 96.8% ($122/126$) recall on this benchmark in the black-box case, and it achieves 100% accuracy in the white-box case. NFARD exhibits different detection results for different types of reuse methods. The greater the modification made by the reuse method to the victim model, the greater the change in neuron functionality, and the more difficult it is for NFARD to detect. For example, quantization and fine-tuning the last layer only slightly modify the model weights, so they are the easiest reuse types to detect and have small distance metrics and large decision values. On the other hand, heterogeneous reuse methods change the model structure, so they are more challenging to detect in the black-box case, indicated by the smaller decision values. Still, NFARD is capable of detecting heterogeneous reuse relations with a success rate of 90% under the black-box setting and 100% under the white-box setting with the help of the linear transformation method. Clearly, surrogate models obtained by different reuse methods all have larger average decision values and smaller average distance metrics compared to negative models, which is in accordance with the conclusion we obtained in the preliminary experiment. Compared with the black-box setting, NFARD performs better in the white-box setting: it achieves 100% precision and 100% recall and yields a greater gap in decision value between the positive and negative suspects. This is as expected since we only have approximated neuron vectors in the black-box case.
V-A2 Comparison with baselines
TABLE III: Comparison with IPGuard, ModelDiff, and DeepJudge in the black-box case. The best F1-score is highlighted.
Reuse type NFARD IPGuard ModelDiff DeepJudge
Fine-tuning - last 12/12 12/12 12/12 12/12
Fine-tuning - all 11/12 12/12 12/12 12/12
Retraining - last 12/12 12/12 12/12 12/12
Retraining - all 12/12 6/12 12/12 12/12
Pruning - 0.3 12/12 12/12 12/12 12/12
Pruning - 0.6 12/12 9/12 12/12 12/12
Quantization - float16 12/12 12/12 12/12 12/12
Quantization - qint8 12/12 12/12 12/12 12/12
Transfer learning 10/12 0/12 11/12 0/12
Knowledge distillation 9/9 0/9 4/9 6/9
Model extraction 8/9 0/9 6/9 6/9
Positive 122/126 87/126 117/126 108/126
Negative 0/60 0/60 28/60 1/60
F1-score 0.984 0.817 0.863 0.919
We compare NFARD with three model reuse detectors that require generating adversarial examples, including IPGuard [20], ModelDiff [25], and DeepJudge [24]. Among them, IPGuard and ModelDiff are black-box detectors, and DeepJudge offers both black-box and white-box versions. Therefore, we compare NFARD with these three methods in the black-box setting and only with DeepJudge in the white-box setting.
The evaluation results under different settings are shown in Table III and Table IV, respectively. We can see that NFARD performs slightly worse than other methods in the black-box case when detecting homogeneous reuse cases, as it fails on one case while other methods all succeed. However, when it comes to heterogeneous reuse cases, NFARD outperforms other methods. In the black-box case, IPGuard and DeepJudge cannot handle transfer learning reuse cases since the classification task is changed. ModelDiff successfully detects one more transfer learning reuse case than NFARD. However, it cannot handle knowledge distillation and model extraction cases as well as NFARD, just like the other two method. As for the white-box setting, DeepJudge cannot handle knowledge distillation and model extraction because of different hidden layer structures. High precision is another advantage of NFARD. The design of the decision criteria of NFARD is more robust than other methods, and therefore NFARD is less likely to produce false alarms. Overall, NFARD has an F1-score of 0.984 in the black-box case and 1.0 in the white-box case, both of which are higher than other DNN reuse detectors.
TABLE IV: Comparison with DeepJudge in the white-box case. The best F1-score is highlighted.
Reuse type NFARD DeepJudge
Homogeneous 96 96
Transfer learning 12/12 12/12
Knowledge distillation 9/9 0/9
Model extraction 9/9 0/9
Positive 126/126 108/126
Negative 0/60 0/60
F1-score 1.0 0.919
Answer to RQ1: NFARD achieves an F1-score of 0.984 in the black-box case and an F1-score of 1.0 in the white-box case on the Reuse Zoo benchmark, outperforming previous methods. Besides, NFARD performs well on various heterogeneous reuse cases, which many baseline methods fail to detect.
V-B RQ2: Application Scope
TABLE V: Qualitative comparison of the application scope of NFARD with adversarial example-based methods.
Factor NFARD IPGuard ModelDiff DeepJudge
Heterogeneous cases Yes No Partial Partial
Fully black-box Yes No No No
Require gradients No Yes Yes Yes
Efficiency High Low Medium Medium
Require labels No Yes No Yes
TABLE VI: Test suite generation time (in seconds) of reuse detectors. The shortest generation time is highlighted.
Detector ResNet-18 ResNet-50 MobileNetV2 GoogLeNet
NFARD 25.2 69.2 27.8 28.5
IPGuard 846.7 4405.2 1831.1 2826.2
ModelDiff 69.5 198.6 80.3 89.1
DeepJudge-black 60.8 152.1 72.4 77.8
DeepJudge-white 141.4 488.5 2064.0 306.4
We have shown that NFARD is capable of detecting more reuse types than other DNN reuse detection methods by employing the linear transformation method. To answer the second research question, we further compare the application scope of NFARD with other methods in the following aspects: model access, scalability, and label information. We provide a brief summary in Table V and describe them in detail.
V-B1 Model access
NFARD requires less model access compared with adversarial example-based DNN detection methods. In the black-box case, NFARD works in a fully black-box manner, while adversarial example-based DNN reuse detectors need gradients of the victim model to generate test samples. For models already deployed on edge devices, even the model owner may not have access to the gradients, which makes the adversarial example-based methods inapplicable. Although ModelDiff also has a fully black-box version, its performance severely degrades, and the authors consider it immature. It is possible to detect DNN reuse relations by generating adversarial examples in a fully black-box manner, but it would lead to a higher cost of the generation process.
V-B2 Scalability
Another important factor that affects the application scope is scalability, which is mainly limited by the efficiency of test suite generation as the metric computation and decision-making process usually takes no more than a few seconds. We report the average test suite generation time for different model architectures of NFARD and other adversarial example-based methods (with $n=1000$) in Table VI. The test suite generation process for NFARD is the same in black-box and white-box cases, while DeepJudge takes two different procedures for test suite generation in each case. NFARD is $2\sim 99$ times faster than the baseline reuse detectors, and its efficiency is also less influenced by the complexity of the model. Since NFARD directly selects samples from the training dataset, the test suite generation time stays almost the same as the number of test cases grows. On the contrary, the test suite generation time of adversarial example-based reuse detectors grows linearly with the number of test cases.
V-B3 Data label
We remark that one of the state-of-the-art methods DeepJudge, requires labeled samples to compute the similarity metrics, and IPGuard needs label information to generate test samples near the decision boundary. This is not an unusual requirement since we also assumed that a subset of training samples is available for detection. However, in some rare situations where normal samples are not labeled, such a requirement may limit the applicability of the detection method. In contrast, NFARD and ModelDiff need no label information throughout the detection process, making them slightly more suitable for rare cases than the above methods.
Answer to RQ2: NFARD requires no gradients of the victim model and no label information from the normal samples, and it is $2\sim 99$ times faster than adversarial example-based methods, thus having better scalability. Therefore, NFARD has a broader scope of application than previous methods.
V-C RQ3: Influencing Factors
Next, we explore how different configurations affect the performance of NFARD. First, we investigate the effectiveness of logarithm approximation in the black-box case by an ablation experiment. Then, we test the performance with different parameters, including decision parameter $\alpha$, target layer choice, and test suite size $n$.
V-C1 Logarithm approximation
TABLE VII: Evaluation results of NFARD without taking the logarithm under the black-box setting.
Reuse type Detected Decision $dist_{eu}^{\prime}$ $dist_{ac}^{\prime}$
Fine-tuning - last 12/12 44.61$\pm$11.0 0.26$\pm$0.15 0.001$\pm$0.0
Fine-tuning - all 10/12 15.08$\pm$13.6 3.04$\pm$1.76 0.23$\pm$0.08
Retraining - last 12/12 32.03$\pm$8.57 1.28$\pm$0.62 0.10$\pm$0.05
Retraining - all 11/12 7.772$\pm$6.78 3.26$\pm$1.67 0.28$\pm$0.07
Pruning - 0.3 12/12 25.76$\pm$12.1 2.13$\pm$1.20 0.14$\pm$0.04
Pruning - 0.6 12/12 22.57$\pm$9.06 2.11$\pm$1.07 0.17$\pm$0.05
Quantization - float16 12/12 44.98$\pm$10.9 0.02$\pm$0.01 0.0$\pm$0.0
Quantization - qint8 12/12 44.87$\pm$11.0 0.09$\pm$0.05 0.0$\pm$0.0
Transfer learning 10/12 0.361$\pm$0.40 0.31$\pm$0.24 0.01$\pm$0.01
Knowledge distillation 6/9 8.158$\pm$14.1 2.59$\pm$1.34 0.22$\pm$0.08
Model extraction 4/9 2.403$\pm$15.4 2.66$\pm$1.34 0.27$\pm$0.10
Positive 113/126
Negative 4/60 -12.57$\pm$6.03 5.21$\pm$2.28 0.44$\pm$0.11
We report the evaluation results of NFARD without taking the logarithm in the black-box setting in Table VII. Clearly, there is a significant drop in performance without the logarithm approximation, especially for knowledge distillation and model extraction reuse cases. The decision value also decreases for most reuse cases, suggesting that the logarithm approximation method does amplify the difference between independently trained models and surrogate models in the black-box case.
V-C2 Decision parameter
(a) Black-box
(b) White-box
Figure 3: ROC curve of NFARD with varying decision parameter $\alpha$ under black-box and white-box settings.
The decision parameter $\alpha$ balances sensitivity and specificity of NFARD. A higher $\alpha$ value yields more conservative detection, reducing false alarms but also recall rates. We show the ROC curve of NFARD with varying $\alpha$ under black-box and white-box settings in Figure 3. The area under the ROC curve for the two cases is 0.998 and 1.0, respectively. Note that the ROC curves are plotted by varying $\alpha$ instead of directly changing the decision thresholds. The evaluation indicates that the performance of NFARD is quite robust to the choice of $\alpha$. We recommend using $\alpha<1$ for the black-box case and $\alpha>2$ for the white-box case in practice.
V-C3 Target layer
We evaluate how the choice of target layer affects the performance of NFARD in the white-box case by running the detection on ResNet-18 victim models with the target layer at different depths. Figure 4 shows that the distance metric gap between surrogate models and independently trained reference models tends to decrease as the target layer goes deeper. This finding is consistent with the suggestion of choosing a shallow layer as the target layer in [24]. There are two possible reasons: one is that shallow layers learn low-level features, which tend to be similar despite being modified by reuse methods; the other is that the model deviation accumulates as the target layer goes deeper.
Figure 4: Distance metrics of different reuse cases under the white-box setting with target layer at different depths.
V-C4 Test suite size
We evaluate NFARD on Reuse Zoo with varying test suite sizes under black-box and white-box settings and report the detector F1-score in Figure 5. We can see that NFARD achieves a perfect 1.0 F1-score under the white-box setting with different test suite sizes, indicating that 100 test samples are sufficient for effective detection in the white-box case. On the other hand, the F1-score curve of NFARD under the black-box setting increases as the test suite size increases until it reaches a plateau with the value of $n$ around 800. In our experiments, we set the test suite size $n$ as 1000 to balance the trade-off between effectiveness and efficiency.
Figure 5: F1-score of NFARD under black-box and white-box settings with different test suite size.
Answer to RQ3: Logarithm approximation contributes to the effectiveness of NFARD in the black-box case. Using a shallow layer as the target layer performs better in the white-box case. NFARD is quite robust to decision parameter $\alpha$ under both circumstances, and we recommend using 1000 test samples in practice.
VI Threats to Validity
Both the proposed method NFARD and the baseline methods have several parameters that affect the performance, and this may affect the validity of the evaluation conclusions if these parameters are not set properly. To mitigate this, on the one hand, we follow the recommended parameter settings by the authors of the baseline methods to ensure their optimal performance. On the other hand, we carry out a series of experiments to explore the influence of NFARD parameters and provide the recommended settings.
The quality of the proposed benchmark is another potential threat. Unbalanced reuse samples, limited kinds of reuse techniques, and unqualified DNN models could put the validity of conclusions at risk. As mitigation, firstly, we include many independently trained models in the benchmark to balance the ratio of positive and negative samples (126: 120). Secondly, we apply seven commonly used reuse techniques with different hyperparameters (e.g., global pruning with varying pruning ratios) to construct the benchmark. Lastly, we also refer to some of PyTorch’s official training recipes [43] to ensure all trained models have qualified performance. In addition, we release the source code for building the benchmark and binary models for future study.
The Reuse Zoo benchmark mainly focuses on image classification tasks and convolutional neural networks, and other learning tasks (e.g., text classification) and model structures (e.g., RNNs and transformers) are not included in our evaluation. Therefore, the conclusions obtained from our experiments may not generalize to other learning tasks and DNN structures. Most existing studies on DNN copyright protection face such risk, except that [24] also evaluated their method on a simple LSTM model trained for an audio classification task. We believe that assessing the performance of copyright protection methods in more diverse scenarios could better demonstrate their usability, and we leave this as future work.
VII Related Work
VII-A DNN Watermarking and Fingerprinting
Watermarking and fingerprinting are two main approaches for passive DNN intellectual property protection [44]. Watermarking methods embed owner-specific information into DNN for IP violation detection. White-box watermarking methods [13, 14, 15] exploit the over-parameterization property of neural networks to plant secret signatures into the model weights through regularization during training. The ownership can be checked by extracting the signatures from the model weights. Black-box watermarking methods [16, 17] use a backdoor attack to implant a watermark in the model during training. The watermark can be extracted by querying the model with triggering samples. Either way, watermarking methods need to be applied in the training process, which will affect the model performance, and there are also methods [18, 19] proposed for removing the watermarks.
Fingerprinting and testing-based similarity comparison methods are post hoc detection approaches that do not tamper with the training process. They extract characterizing features (sometimes called fingerprints) from the models using carefully generated input samples close to the decision boundary [20] or conferrable adversarial examples [21]. A recent study [22] utilizes universal adversarial perturbations [45] to achieve similar effects. ModelDiff [25] and DeepJudge [24] are the two state-of-the-art testing-based methods. ModelDiff can handle the transfer learning reuse case, and DeepJudge has a white-box detection procedure and achieves a higher detection accuracy on homogeneous reuse cases.
VII-B Active DNN Copyright Protection
Different from passive protection methods, active protection aims to prevent copyright infringement from occurring. For distributed model copies, DNN authentication techniques [46, 47, 48] manage the authorized usage of the source model by encoding the model with an authentication key, so that only authorized users with valid keys can access the model normally, while adversaries with invalid keys can only use the model with greatly reduced performance, thus preventing unauthorized model usage. As for MLaaS mode, model extraction is the major threat, and inference perturbation techniques [49, 50, 51] are proposed to defend the source model by modifying the inference results before returning them to the users.
VII-C Neural Network Representation
Our approach is inspired by the line of research on neural network representation. Canonical correlation analysis techniques such as SVCCA [26] and PWCCA [27] are proposed to analyze the representation of deep neural networks. These analysis techniques contribute to understanding the interpretability of neural networks and learning dynamics. Previous studies concentrate on revealing the individual differences among independently trained models [28] and the properties of converged weights of neural networks [27] by extensive experiments. The results indicate that independently trained neural networks usually have divergent neuron functionalities, even though they have the same architecture and are trained with the same dataset. By exploiting the fact that models obtained by reuse techniques have similar neuron functionalities, we devise our adversarial example-free detection method.
VIII Conclusion
In this paper, we propose NFARD to protect deep neural network copyright based on measuring the neuron functionality similarity of two models. We devise a linear transformation method which enables NFARD to detect heterogeneous reuse cases that existing approaches could not handle well. To better evaluate copyright protection methods, we construct a reuse detection benchmark covering various commonly used reuse methods and popular model architectures. Empirical evaluation results show the effectiveness and efficiency of our approach in both black-box and white-box settings, outperforming existing methods. In the future, we will explore the application of our approach to other learning tasks and model structures, e.g., language models for text classification.